supplies: $1,700
operating expenses: $9,200
Accounts payable: $10,000
Accounts Receivable: $3,000
Common Stock:$9,000
Retained Earnings (begining): $5,700
Service Revenue: $19,000
Cash: $14,000
Dividends: $6,500
Notes Payable: $2,000
Equipment: $11,300
Calculate the following: Net Income/(Net Loss), Ending retained Earnings, and total assets?
Answer #1

Net income= Service revenue- Operating expenses = 19,000- 9200= 10,800

Ending retained earnings= Beginning retained earnings + Net income - Divided = 5700+ 10800- 6500= 10,000

Total assets= account receivable + cash +supplies + equipment

Total assets = 3000+ 14000+ 1700 +11300= 30,000